OpenStudy (anonymous):

Evaluate the integral (2x +3)/(x+1)^2 dx from 0 to 1 with integration by parts.

OpenStudy (anonymous):

\[\frac{2x+3}{(x+1)^2}=\frac{x+1}{(x+1)^2}+\frac{x+3}{(x+1)^2}\] might be a good first step

OpenStudy (carson889):

Should it not be: \[\frac{ 2x }{ (x+1)^{2} } + \frac{ 3 }{ (x+1)^{2} }\]

OpenStudy (anonymous):

that would work too

OpenStudy (raden):

alternative u can use int by u-sub let u = x+1 --------> x = u - 1 du = dx see ur integral becomes int (2(u-1)+3)/u^2 du = int (2u + 1)/u^2 du = int (2/u + u^-2) du (i miss the limits for while)

OpenStudy (raden):

that should be easier for u, @Prof :)

OpenStudy (raden):

btw, i didnt read above u want solving by parts sorry :)